1. Calculate the turning points of the function y=sin3t using differential calculus
  2. Show which are maxima, minima or points of inflexion using the second derivative
1
Expert's answer
2021-11-04T09:19:44-0400

Domain of the function is not given, so let us take

"t\in (0,2)"

"y=sin3t"

To find the maximum and minima, we need to find the "y\prime" and set this to zero.


"y = sin 3t"

differentiate with respect to "t"


"y\prime = 3 cos 3t = 0""3t = \frac {\pi}{2}, \frac {3\pi}{2}""t = \frac {\pi}{6}; \frac {\pi}{2} = 0.52; 1.57"

Now,


"y\prime\prime = - 9 sin 3t"

"y\prime\prime (0.52)=-9sin3(0.52)=-9sin(1.56)=-8.999
\\y\prime\prime (0.52) < 0\\
So, y \space has \space maximum \space at \space t = 0.52"

Again,

"y\prime\prime (1.57)=-9sin3(1.57)=-9sin(4.71)=8.999
\\y\prime\prime (1.57) > 0\\
So, y \space has \space minimum \space at \space t = 1.57"

To find the Inflection point set "y\prime\prime = 0"

"- 9 sin 3t = 0\\
sin 3t = 0 \space at \space 3t = \pi" or 3.14

inflection point is at "t = 1.04"

Now plotting the graph of function to show the turning point , maxima ,minima



Turning point of function is t = 1.04 for "t\in (0,2)" . Turning point of function is the point where concavity of function is changes. It is also called point of inflection.